The table below shows the parameters for the economy of Hutu. C = 70 + 0.6Y XN = 150 − 0.1Y I = 140 G = 240

The value of equilibrium income is $ 1200 .

d.if full employment income is $925, what change in government spending is necessary to move the economy to this level? Government spending needs to by $
Answer #1

Let new Government Spending level = G`

Aggregate demand AD = C + G`+ I + NX

AD = 70+.6Y +140 + G` +150-.1Y

AD = 360+ .5Y+G`

Now at eqm, Y = AD

so, Y = 360+ .5Y + G`

.5 Y = 360 + G`

put Y = 925

.5*925 = 360 + G`

G` = 102.5

so ∆G = 102.5 - 240

= -137.5

​​​​​​decrease the G by 137.5